OpenStudy (anonymous):

Calculate the mass of methane that must be burned to provide enough heat to convert 113.0 g of water at 19.0°C into steam at 125.0°C. (Assume that the H2O produced in the combustion reaction is steam rather than liquid water.)

OpenStudy (zpupster):

Temperature difference is (125 - 19) = 106 C. Energy needed to raise the temperature of 113 gms by 106 degrees will be Q = m*s*t Q ---is heat energy m---- is the mass of water s---is the specific heat (of H2O 4.2 J/g˚C) t---- is the temperature difference. => Q = 119 * 4.2 * 106 => Q = 52979 Now methane liberates 55 kJ of heat energy per gram. So, methane required will be 52979/55 = 963.3 grams